39.0 mL of a pure, water-soluble liquid alcohol with a density of 1.854 g/mL is diluted with water, and the final mass of the solution is 5.00 × 10g. What is the mass percent of the alcohol in this solution?


1
Expert's answer
2021-05-17T04:07:00-0400

mass = density × volume

mass of alcohol in the solution = 39.0 mL × 1.854 g/mL = 72.31g


final mass of the solution = 5.00 × 10² g = 500g


mass percent of the alcohol = mass of alcohol/total mass of solution × 100% = 72.31/500 × 100% = 14.462%


"\therefore" The mass percent of the alcohol in this solution is "14.462\%"